OpenStudy (anonymous):

Medal and fan:) 5.) The ordered pairs (1,2), (2,4), (3,8), (4,16), and (5,32) represents a function. What is a rule that represents this function?

OpenStudy (anonymous):

A. y=x^2 B.y=2x c.y=2^x D.y=x+2

OpenStudy (anonymous):

C , 2^X

OpenStudy (demonchild99):

its just timsing everything by two

OpenStudy (anonymous):

Thank you i got stuck on this one lol

OpenStudy (anonymous):

i also have another one like this

OpenStudy (demonchild99):

lol....i cant do anything with graphs...lol....put it up

OpenStudy (anonymous):

demon, the Y- value relates to each other by times-ing 2 , but the relation of X and Y is that Y= 2^X

OpenStudy (demonchild99):

yeah

OpenStudy (anonymous):

The ordered pairs (1,36), (2,49), (3,64), (4,81) and (5,100) A.y=x^2 B.36x C.y=(x+5)^2 D.y=(x+6)^2

OpenStudy (anonymous):

Dontspeaktome, using my explanation to demon, could you figure it out?

OpenStudy (anonymous):

Try to see how the (X,Y) X relate to Y, sub the values into the equation and see if u get the function values. it wud be C.

OpenStudy (anonymous):

Oh ok. i was trying to figure it out but i couldnt so thank you so much